### 06 Mar 2026 v1

# The Coordination Gap: Alternation Metrics for Temporal Dynamics in Multi-Agent Battle of the Exes

Abstract
Multi-agent coordination dilemmas expose a fundamental tension between individual optimization and collective welfare, yet characterizing such coordination requires metrics sensitive to temporal structure and collective dynamics. As a diagnostic testbed, we study a BoE-derived multi-agent variant of the Battle of the Exes, formalizing it as a Markov game in which turn-taking emerges as a periodic coordination regime. Conventional outcome-based metrics (e.g., efficiency and min/max fairness) are temporally blind—they cannot distinguish structured alternation from monopolistic or random access patterns—and fairness ratios lose discriminative power as $n$ grows, obscuring inequities.
To address this limitation, we introduce Perfect Alternation (PA) as a reference coordination regime and propose six novel Alternation (ALT) metrics designed as temporally sensitive observables of coordination quality. Using Q-learning agents as a minimal adaptive diagnostic baseline, and comparing against random-policy null processes, we uncover a clear measurement failure: despite exhibiting deceptively high traditional metrics (e.g., reward fairness often exceeding $0.9$), learned policies perform up to $81\%$ below random baselines under ALT-variant evaluation—a deficit already present in the two-agent case and intensifying as $n$ grows.
These results demonstrate, in this setting, that high aggregate payoffs can coexist with poor temporal coordination, and that conventional metrics may severely mischaracterize emergent dynamics. Our findings underscore the necessity of temporally aware observables for analyzing coordination in multi-agent games and highlight random-policy baselines as essential null processes for interpreting coordination outcomes relative to chance-level behavior.
keywords:
Fair Division , Multi-Agent Systems , Welfare Measurement , Turn-Taking , Coordination Dynamics , Markov Games , Cooperative Games
journal: Mathematical Social Sciencesorcid1orcid1footnotetext: ORCID: https://orcid.org/0000-0003-1842-8227orcid2orcid2footnotetext: ORCID: https://orcid.org/0000-0003-0020-6394
1 Introduction
Coordination among multiple self-interested agents is a fundamental problem in complex systems, where local decision rules and asymmetric incentives give rise to emergent collective behavior Rankin et al. (2007); Axelrod (1984); Perolat et al. (2017). In the absence of centralized control, such systems exhibit interdependent feedback between individual actions and collective outcomes, producing diverse long-run coordination regimes Vainer and Kukacka (2021); Maruta and Okada (2012); Zhu et al. (2020); Shan et al. (2019); He et al. (2015); Tian et al. (2026); Pikovsky (2009). A central challenge is to characterize how efficient or equitable collective patterns emerge—and persist—purely through repeated interaction, adaptation, and feedback. Such coordination problems arise across diverse domains and turn-taking phenomena: decentralized access in wireless networks Shenoy (2007), conversational turn-taking in linguistics Sacks et al. (1974), animal communication and rhythmic alternation Takahashi et al. (2013); Greenfield et al. (2021); Ravignani et al. (2019); Pika et al. (2018), stochastic phenotype switching Rainey et al. (2011), vision-based robotic coordination Mezey et al. (2025), and emergent conversational norms in virtual environments Bonfert et al. (2022). Yet existing evaluation frameworks—including welfare measures from social choice theory and allocation criteria from fair division—remain largely insensitive to the temporal structure inherent in these coordination processes.
Repeated interaction among egoistic agents can be naturally viewed as a discrete-time, stochastic, and history-dependent dynamical system Pikovsky (2009); Axelrod (1984). Early computational studies, most notably Axelrod’s investigation of cooperation among egoists without central authority Axelrod (1984), demonstrated that simple adaptive rules can generate stable behavioral regularities through repetition alone. From a game-theoretic perspective, learning rules induce adaptive trajectories in a high-dimensional strategy space, where long-run behavior corresponds to recurrent coordination regimes rather than one-shot equilibria Pikovsky (2009). Understanding which coordination patterns emerge, and how to distinguish them quantitatively, remains an open problem.
The Battle of the Exes (BoE) game provides a minimal yet expressive model for studying such coordination dynamics Hawkins and Goldstone (2016). Unlike cooperative dilemmas, BoE is a coordination game with asymmetric payoffs, where agents must resolve conflict over preferred outcomes. In repeated play, the socially optimal solution is not static cooperation but temporal alternation, in which agents take turns accessing the higher-reward outcome Hawkins and Goldstone (2016); Papadopoulos and Sanchez-Fibla (2021); Papadopoulos (2020). This alternation corresponds to a structured temporal pattern—analogous to a periodic orbit—rather than a simple payoff equilibrium. As such, BoE offers a natural testbed for examining the emergence of turn-taking conventions as collective dynamical patterns.
Despite extensive prior work on the repeated BoE, existing studies are almost exclusively restricted to the two-agent case Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023); Lopez Hernandez (2022). This limitation obscures coordination phenomena that arise in more realistic multi-agent settings, where competition among several agents induces richer and higher-dimensional dynamics. Moreover, considering BoE-derived multi-agent variants exposes a critical methodological gap: commonly used fairness and efficiency metrics are typically time-averaged and payoff-based Gini (1912); Theil (1967); Allison (1978); Jiang and Lu (2019); Perolat et al. (2017), rendering them insensitive to temporal ordering. As a result, such metrics cannot distinguish genuine turn-taking from monopolistic, random, or weakly correlated access patterns, particularly in multi-agent scenarios where high aggregate efficiency may coexist with poor coordination.
In this study, we address these limitations by formalizing a BoE-derived multi-agent dynamic Battle of the Exes (MBoE) as a Markov game Littman (1994). To characterize turn-taking as a temporal phenomenon, we introduce the concept of a Perfect Alternation (PA) reference coordination regime for idealized turn-taking. Building on this benchmark, we propose six novel Alternation (ALT) metrics designed as time-series observables that explicitly capture temporal structure and coordination quality Papadopoulos and Sanchez-Fibla (2021); Papadopoulos et al. (2025). Through extensive numerical simulations with Q-learning agents—interpreted here as a discrete stochastic adaptive process and used as a minimal learning baseline—we demonstrate that traditional metrics can mask coordination failure, while ALT metrics reliably distinguish qualitatively different long-run regimes. Our results highlight the importance of temporally sensitive observables for analyzing coordination dynamics in strategic multi-agent interaction.
1.1 Main Contributions
The main contributions of this study are as follows:
  1. 1.
    Multi-agent formalization of BoE: We study a BoE-derived multi-agent variant of the classical two-player Battle of the Exes and formalize it as a Markov game, enabling the study of high-dimensional coordination dynamics.
  2. 2.
    Perfect Alternation (PA): We introduce the concept of Perfect Alternation (PA) as a reference coordination regime for turn-taking tasks in multi-agent systems Papadopoulos (2020); Papadopoulos and Sanchez-Fibla (2021).
  3. 3.
    Novel Alternation (ALT) metrics: We propose six ALT measures that overcome the temporal insensitivity and coordination blindness of traditional fairness and efficiency metrics—their inherent inability to distinguish structured alternation from monopolistic or random access patterns, even when reporting values approaching 1.0. This dual limitation becomes severe and misleading as the number of agents grows, with traditional metrics masking severe coordination failure Papadopoulos and Sanchez-Fibla (2021); Papadopoulos et al. (2025).
  4. 4.
    ALT ratio benchmarking framework: We develop a regression-based methodology mapping ALT metric values to Perfect Alternation–equivalent performance, enabling interpretable quantification of coordination quality.
  5. 5.
    Random baseline as null process: We establish random policy baselines as a statistical null hypothesis for coordination, demonstrating that high fairness or efficiency values can arise purely by chance—a methodological practice not consistently reported in prior BoE and closely related coordination studies Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023).
  6. 6.
    Numerical insights into emergent coordination: Through extensive Q-learning simulations, we show that efficient and seemingly fair outcomes may conceal poor coordination, while ALT metrics expose underlying dynamical regimes.
1.2 Research Questions
This study focuses on the following research questions:
  1. RQ1:
    How do conventional fairness and efficiency metrics behave in multi-agent turn-taking settings, and under what conditions do they fail to distinguish structured coordination from random or monopolistic dynamics?
  2. RQ2:
    How effectively do ALT metrics separate coordination regimes in empirical evaluation (perfect alternation, partial alternation, random access) in multi-agent BoE?
  3. RQ3:
    How does coordination quality, as measured by ALT metrics, scale with the number of agents in the multi-agent Battle of the Exes?
  4. RQ4:
    How do ALT metrics compare against a random-policy null process, and what does this comparison reveal about apparent coordination measured by conventional metrics?
Note: Simple Q-learning is used as a minimal adaptive baseline to generate coordination dynamics, rather than as a benchmark for optimal learning performance.
1.3 Manuscript Organization
The remainder of the manuscript is organized as follows. Section 2 reviews the classical Battle of the Exes and introduces the proposed multi-agent variant. Section 3 discusses traditional fairness and efficiency metrics and motivates the need for temporally sensitive observables. Section 4 introduces the Perfect Alternation reference coordination regime and the six ALT metrics. Section 5 presents the numerical experiments and analyzes emergent coordination regimes. Section 6 discusses the findings and their implications. Section 7 concludes with key insights, limitations, and directions for future work.
2 Battle of the Exes and Prior Work
2.1 The Battle of the Exes Game
The Battle of the Exes (BoE), introduced by Hawkins and Goldstone (2016), is a coordination game derived from the classical Battle of Sexes, but with inverted incentives. Rather than coordinating on the same action, agents must avoid each other while competing for asymmetric rewards. If both agents select the same action, no reward is obtained; if they select different actions, each agent receives a payoff determined by individual preference asymmetry.
Table 1: Two-player Battle of the Exes payoff matrix
Action 1 Action 2
Action 1 (0, 0) (3, 2)
Action 2 (2, 3) (0, 0)
BoE exists in two variants: a ballistic and a dynamic formulation. In the ballistic version, agents make simultaneous discrete action choices at each round, analogous to a standard normal-form game. In contrast, the dynamic version embeds decision-making in continuous space and time: agents move toward preferred destinations within an environment, and conflicts arise when agents arrive at the same destination within overlapping time windows. This spatiotemporal coupling introduces additional feedback channels that facilitate the emergence of turn-taking behavior in repeated interaction.
BoE admits multiple equilibria in the one-shot game, including pure and mixed Nash equilibria. However, in repeated interaction the socially optimal outcome is not static coordination but temporal alternation, where agents take turns accessing the higher-reward outcome. This alternation corresponds to a structured temporal pattern, rather than a stationary payoff equilibrium.
Hawkins and Goldstone investigated both ballistic and dynamic versions of BoE using human participants. Their experiments demonstrated that the dynamic formulation substantially increases the emergence of turn-taking behavior, with some pairs converging to near-perfect alternation, highlighting the role of spatiotemporal interaction in coordination dynamics Hawkins and Goldstone (2016).
2.2 Learning-Based Studies of BoE
Following these findings, several studies examined whether artificial agents could reproduce the alternation behavior observed in humans. Puig Camps (2018) investigated repeated BoE using independent Q-learning agents and reported convergence to stable deterministic strategies under specific learning-rate and exploration settings. Gasparrini and Sanchez-Fibla (2018) incorporated loss aversion into agents’ utility functions, accelerating convergence in the ballistic formulation and enabling alternation in the dynamic variant. Related modeling work explored control-based or episodic-memory agents in BoE-like settings, focusing on matching human efficiency/fairness/stability profiles rather than multi-agent scaling Freire et al. (2020, 2023); Lopez Hernandez (2022).
Despite these advances, learning-based BoE studies remain almost exclusively restricted to the two-agent setting, where alternation reduces to binary role switching and random policies can achieve near-perfect Fairness, Reward Fairness, and TT Fairness, and high efficiency by chance alone. Evaluation relies primarily on time-averaged, payoff-based metrics such as efficiency and fairness, which suffer from temporal insensitivity and pattern ambiguity (random access can yield values indistinguishable from coordination) even in dyadic settings, with additional limitations—coordination blindness and exclusion of intermediate agents—emerging as $n$ grows. Moreover, random-policy baselines are not consistently reported as explicit null processes Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023), making it difficult to assess whether observed coordination reliably exceeds chance-level behavior. An exception is recent MSc thesis work that includes randomized control comparisons in dyadic settings Lopez Hernandez (2022); however, this comparison served primarily to validate model performance against human data rather than to establish a formal null hypothesis for multi-agent coordination assessment, leaving open the question of how severely traditional metrics can obscure coordination failure. Our study addresses these gaps by (i) formalizing random policies as an explicit statistical null process with quantitative Coordination Score and Relative Change methodology, (ii) scaling baselines to multi-agent settings ($n\in\{2,3,5,8,10\}$) across four experimental configurations, four outcome-based metrics, and six ALT variants, and (iii) demonstrating that traditional metrics can appear deceptively high even under random play—a finding that motivates the introduction of temporally sensitive ALT metrics. Building upon and substantially extending our early conceptual work on alternation-sensitive metrics and periodicity-based analysis Papadopoulos (2020); Papadopoulos and Sanchez-Fibla (2021); Papadopoulos et al. (2025), this paper introduces a rigorous theoretical framing, calibrated random baselines, and scaled multi-agent evaluation with refined PA and metric definitions.
2.3 Why Multi-Agent BoE Is Fundamentally Different
BoE-derived multi-agent variants introduce qualitatively new coordination challenges. With $n>2$ agents competing simultaneously for access to a single high-reward outcome, alternation is no longer binary but requires collective periodic rotation across all participants. Partial exclusion, intermittent monopolization, and irregular access patterns emerge naturally, producing richer temporal dynamics absent from the two-agent case.
Crucially, this extension exposes a methodological gap: conventional fairness and efficiency metrics scale poorly with agent count and remain insensitive to temporal ordering. High efficiency may result from persistent monopolization or random access, while fairness ratios often collapse to extreme values that obscure intermediate agents. Consequently, even severe coordination failures can appear successful under traditional evaluation.
2.4 Battle of the Exes as a Minimal Multi-Agent MARL Testbed
To address these limitations, we formalize a Multi-Agent Battle of the Exes (MBoE) as a BoE-derived multi-agent variant in the form of an episodic Markov game involving $n\geq 2$ self-interested agents. In each episode, agents act simultaneously and compete to reach a terminal state associated with a shared high reward. If exactly one agent succeeds, it receives the maximum payoff; if multiple but not all agents succeed, they receive reduced rewards; if all agents tie, no reward is assigned. We use a minimal learning setting (tabular Q-learning), a minimal dynamic formulation (single pre-terminal position), and a minimal environment design to maximize interpretability.
This BoE-derived formulation preserves the core incentive structure of BoE—mutual interference with asymmetric rewards—while providing a plausible multi-agent variant. The optimal collective behavior is no longer simple avoidance but periodic turn-taking across all agents. Whether and how such behavior emerges under decentralized learning—and how it should be evaluated—are the central questions addressed in the remainder of this manuscript.
In the following sections, we demonstrate that existing fairness and efficiency metrics are fundamentally inadequate for characterizing multi-agent turn-taking. We therefore introduce Perfect Alternation as a reference coordination regime and propose novel, temporally sensitive Alternation (ALT) metrics capable of distinguishing genuine coordination from spurious access patterns.
2.5 Multi-Agent Battle of the Exes as a Markov Game
We now formally define the Multi-Agent Battle of the Exes (MBoE) as an episodic Markov game Littman (1994). This BoE-derived formulation constitutes a core contribution of this work, providing a plausible multi-agent variant of the classical two-agent coordination tension—mutual interference with asymmetric payoffs.
2.5.1 Game Definition
The MBoE is defined as a Markov game
$\mathcal{G}=(N,S,A,T,R),$ (1)
where $N$ denotes the set of agents, $S$ the state space, $A$ the joint action space, $T$ the transition function, and $R$ the reward function.
The game is episodic: each episode begins from a fixed initial configuration and terminates when at least one agent reaches a terminal state.
2.5.2 Agents
Let
$N=\{1,2,\dots,n\},\quad n\geq 2,$ (2)
denote the set of self-interested agents. All agents act simultaneously and independently at each decision step and aim to maximize their individual cumulative reward.
2.5.3 Action Space
At each time step $t$, each agent $i\in N$ selects an action
$a_{t}^{i}\in\mathcal{A}=\{\text{move},\text{stay}\}.$ (3)
The joint action is denoted by $\mathbf{a}_{t}=(a_{t}^{1},\dots,a_{t}^{n})$.
2.5.4 Transition Dynamics
Agents progress along individual paths of finite length toward terminal outcomes. The transition function
$T:S\times A\rightarrow S$ (4)
is deterministic and updates each agent’s position according to
$p_{t+1}^{i}=\begin{cases}p_{t}^{i}+1,&\text{if }a_{t}^{i}=\text{move},\\ p_{t}^{i},&\text{if }a_{t}^{i}=\text{stay}.\end{cases}$ (5)
Positions are non-decreasing, and agents may advance by at most one position per time step. An episode terminates as soon as at least one agent reaches the terminal position.
2.5.5 Reward Function
Rewards are assigned only at terminal states. Let
$k=\sum_{i=1}^{n}\mathbb{I}(p_{T}^{i}=\text{terminal})$ (6)
denote the number of agents reaching the terminal position at the end of an episode. The reward for agent $i$ is given by
[…]
If all agents reach the terminal position simultaneously ($k=n$), no reward is assigned. This reward structure is inspired by the canonical two-agent Battle of the Exes, preserving its core coordination tension: agents benefit from exclusive access but suffer when competing for the same outcome. While the canonical two-agent BoE is defined over asymmetric destination preferences, the present formulation abstracts this structure into a single contested terminal with exclusivity-based rewards. This abstraction preserves the essential coordination tension—mutual interference under asymmetric benefit—while enabling systematic scaling to the multi-agent regime.
For instance, in a 3-agent system with $r_{\mathrm{high}}=100$: if agent 1 wins alone, $R_{1}=100$ and $R_{2}=R_{3}=0$; if agents 1 and 2 both reach the terminal (partial tie), $R_{1}=R_{2}=r_{\mathrm{low}}$ and $R_{3}=0$; if all three agents tie, $R_{1}=R_{2}=R_{3}=0$.
2.5.6 Coordination Structure
In the multi-agent setting, optimal collective behavior corresponds to periodic turn-taking, where agents alternate access to the high-reward outcome across episodes. This structured temporal pattern forms the basis for the Perfect Alternation reference coordination regime introduced in Section 4.
3 Traditional Multi-Agent Fairness Measures
Multi-agent fairness research distinguishes between outcome fairness (cumulative payoff distribution) and procedural fairness (temporal distribution of rewards). Common outcome measures include the Gini coefficient Gini (1912), Theil index Theil (1967), and Coefficient of Variation Allison (1978), while recent multi-agent work often uses payoff-based fairness or equality variants Jiang and Lu (2019); Perolat et al. (2017). In the fair division literature, allocation mechanisms are evaluated by properties such as envy-freeness, proportionality, and strategyproofness Suksompong and Teh (2023); Camacho et al. (2023), yet these criteria remain distributional and do not account for the temporal sequence of allocations. Similarly, axiomatic approaches to cooperative surplus sharing emphasize distributional fairness and consistency Calleja and Llerena (2020), but do not address the temporal ordering of payoffs. Consequently, these metrics fail to capture when agents receive rewards—critical for turn-taking tasks where an agent monopolizing early episodes violates fairness even if cumulative payoffs equalize. These shortcomings already appear in the two-agent BoE setting; see Supplementary Section S3 for extended discussion. This limitation motivates the alternation-sensitive measures developed in Section 4.
3.1 Limitations of Traditional Metrics in BoE and MBoE
We examine four traditional metrics and their shortcomings for the Battle of the Exes (BoE) and its multi-agent variant (MBoE).
Fairness.
Defined as $F^{\nu}=\min_{i}w_{i}/\max_{i}w_{i}$. Key limitations: (1) ignores intermediate agents (only min/max), (2) undefined when all episodes are ties ($0/0$), and (3) identical values arise from fair rotation, subgroup monopoly, or random access.
Efficiency.
Defined as $E^{\nu}=\sum_{i}r_{i}^{\text{total}}/(\nu\times r_{\text{high}})$, measuring total reward capture. Key limitations: (1) high efficiency can result from monopoly or alternation, and (2) provides no information about coordination patterns.
Turn-Taking Fairness.
Defined as $F_{\nu}^{TT}=\min_{i}t_{i}/\max_{i}t_{i}$ Hawkins and Goldstone (2016). Key limitations: (1) still ignores intermediate agents, and (2) high $t_{i}$ can reflect frequent collisions rather than genuine alternation.
Reward Fairness.
Defined as $F_{\nu}^{R}=\min_{i}p_{i}/\max_{i}p_{i}$ Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018), comparing cumulative payoffs. Key limitations: (1) high values can result from tie rewards without turn-taking, and (2) a dominant subgroup can still yield high Reward Fairness.
Here, $w_{i}$ counts exclusive wins, $t_{i}$ counts terminal arrivals (including ties), and $p_{i}$ is cumulative payoff for agent $i$ over $\nu$ episodes.
Table 2: Diagnostic limitations of traditional metric combinations. Each pairing was systematically evaluated across all nine value combinations arising from Efficiency $\in\{0,(0,1),1\}$ and the respective Fairness variant $\in\{0,(0,1),1\}$. Temporal coordination patterns remain invisible in all cases; fair outcome interpretation is possible only for Reward Fairness at the optimal corner (Efficiency = 1, Reward Fairness = 1). Complete cell-by-cell analysis appears in Supplementary Material §3.2.5.
Metric Pair Temporal Fair Outcome
(with Efficiency) Structure Confirmed?
Fairness None No (ignores intermediate agents)
Turn-Taking Fairness None No (conflates collisions with turns)
Reward Fairness None Only at (Eff=1, RF=1)
Combining Efficiency with Fairness or TT Fairness does not even guarantee a fair outcome interpretation: Fairness ignores intermediate agents, while TT Fairness conflates exclusive wins with collision-induced arrivals. Only Reward Fairness, when combined with high Efficiency, can indicate equitable payoff distribution—but even then, temporal structure remains invisible. This double inadequacy—failing both as fairness indicators and as coordination measures—motivates the development of ALT metrics. Table 2 illustrates that interpretable patterns emerge only at extremes; at intermediate values—common in practice—these metrics fail to distinguish coordination quality. Random-policy baselines are not consistently reported as explicit null references in prior BoE-related studies Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023); we therefore include them as an explicit null reference for coordination assessment.
4 Proposed Assessment Framework
Traditional metrics fail to capture temporal structure—how access to high-payoff states is distributed across agents over time. We introduce Alternation Measures (ALT) that explicitly quantify who succeeds, when, and in what sequence.
4.1 Perfect Alternation as Reference
Perfect Alternation (PA) is a coordination regime where each of $n$ agents reaches the highest-payoff state exactly once within every block of $n$ episodes, in any order. Stricter order-sensitive variants (e.g., requiring each agent to occupy each rank equally often across batches) are possible but not imposed here. PA can be used descriptively, normatively, or prescriptively depending on the application; here it serves only as a reference regime against which observed coordination can be compared Papadopoulos (2020); Papadopoulos and Sanchez-Fibla (2021). Human participants in BoE experiments sometimes converge to PA-like patterns Hawkins and Goldstone (2016), motivating its use as a meaningful boundary condition. PA is Pareto-optimal under strict rotation, and in the two-agent ballistic case it constitutes a Nash equilibrium; detailed equilibrium links and the three motivating reasons (empirical observability, temporal sensitivity, metric calibration) are provided in Supplementary Section S4.
4.2 ALT Framework
Given $\nu$ episodes and $n$ agents, ALT metrics evaluate all $b=\nu-(n-1)$ overlapping batches of $n$ consecutive episodes. Each batch $j$ receives a weight $\beta_{j}\in[0,1]$, where $\beta_{j}=1$ indicates perfect alternation. The overall score averages batch weights:
$\mathrm{ALT}=\frac{1}{b}\sum_{j=1}^{b}\beta_{j}$ (8)
ALT measures satisfy: (i) boundedness in $[0,1]$; (ii) $\mathrm{ALT}=1$ under PA; (iii) monotonic degradation under monopolization or ties; (iv) sensitivity to exclusivity; (v) permutation invariance Papadopoulos and Sanchez-Fibla (2021); Papadopoulos et al. (2025). For this study, order within a batch is not penalized; any permutation is treated as ideal. Full symbol definitions and computation details are provided in Supplementary Section S4.
4.3 Six ALT Variants
We define six variants spanning tolerant to strict formulations. Let $f_{j}$ = unique winners in batch $j$, $\tau_{j}$ = total terminal occurrences, $w_{j}$ = episodes with exclusive winners, $g_{j}$ = agents with exactly one exclusive win, and $Y_{\ell}$ = agents reaching terminal in episode $\ell$.
FALT (Fractional).
Most tolerant; measures unique winners relative to terminal occurrences:
$\beta_{j}^{\mathrm{FALT}}=\frac{f_{j}}{\tau_{j}}$ (9)
qFALT (Quadratic Fractional).
More sensitive (quadratic penalty): [^1] Previously called EFALT; renamed since the operation is quadratic, not exponential.
$\beta_{j}^{\mathrm{qFALT}}=\left(\beta_{j}^{\mathrm{FALT}}\right)^{2}$ (10)
EALT (Exclusive).
Emphasizes exclusivity—episodes with single winners:
$\beta_{j}^{\mathrm{EALT}}=\frac{w_{j}\cdot f_{j}}{n^{2}}$ (11)
qEALT (Quadratic Exclusive).
More sensitive exclusive measure (quadratic penalty): [^2] Previously called EEALT.
$\beta_{j}^{\mathrm{qEALT}}=\left(\beta_{j}^{\mathrm{EALT}}\right)^{2}$ (12)
CALT (Complete).
Primary metric; penalizes ties explicitly:
$\beta_{j}^{\mathrm{CALT}}=\frac{\sum_{\ell=1}^{n}(n-Y_{\ell})\cdot\beta_{j}^{\mathrm{qFALT}}}{n(n-1)}$ (13)
AALT (Absolute).
One of the strictest; rewards only agents with exactly one exclusive win per batch:
$\beta_{j}^{\mathrm{AALT}}=\frac{g_{j}}{\tau_{j}}$ (14)
CALT serves as the primary metric due to its balanced treatment of exclusivity and ties. FALT provides a tolerant baseline, EALT isolates exclusivity, and AALT sets a strict diagnostic bound. ALT measures complement—rather than replace—traditional metrics by capturing temporal coordination invisible to outcome-based evaluation.
5 Experimental Setup and Results
5.1 Environment and Experimental Design
We study MBoE coordination using two state representations: Type-A (position-only: $s_{t}^{A}=(p_{t}^{1},\dots,p_{t}^{n})$) and Type-B (with episodic memory: $s_{t}^{B}=((p_{t}^{1},\dots,p_{t}^{n}),(z_{t-1}^{1},\dots,z_{t-1}^{n}))$, where $z_{t-1}^{i}=1$ if agent $i$ won the previous episode). Two reward schemes scale tie penalties: ILF ($r_{\mathrm{low}}=r_{\mathrm{high}}/n$) and IQF ($r_{\mathrm{low}}=r_{\mathrm{high}}/n^{2}$), with $r_{\mathrm{high}}=100$. [^3] Previously called EFR/EAFR; renamed to reflect inverse polynomial operations.
Experiments span all combinations of state types (Type-A/B), reward schemes (ILF/IQF), and agent counts ($n\in\{2,3,5,8,10\}$), yielding 20 configurations. Code is available at https://github.com/dentros/Alternation (GNU GPL).
5.2 Q-Learning and Training
Each agent maintains an independent tabular Q-table with no inter-agent communication. Hyperparameters: $\gamma=0.999$, $\alpha=0.3$, $\epsilon$-greedy with linear decay from $\epsilon_{initial}=0.9$ to $\epsilon_{min}=0.004$. Q-learning was selected over deep RL for complete interpretability of coordination dynamics.
Episode counts scale with agent complexity:
$\text{Episodes}(n)=\text{BASE}\times\left(\frac{n}{2}\right)^{2}\times\left(1+\ln\left(\frac{n!}{2!}\right)\right)$ (15)
with BASE=1000, yielding: 1,000 (2 agents), 4,721 (3), 31,839 (5), 174,583 (8), and 385,281 (10 agents) episodes per configuration. This schedule is heuristic, chosen to balance state-space growth against computational and energy constraints; pilot runs with alternative scalings produced qualitatively similar trends.
5.3 AltRatio and PA-Equivalent Interpretation
To interpret ALT values, we define $\text{AltRatio}=x/n\in[0,1]$, where $x$ represents the equivalent number of perfectly alternating agents, while the remaining nxn-x agents stay stationary. Multiplying by $n$ yields the PA-equivalent: “This system coordinates as well as $x$ out of $n$ perfectly alternating agents.” Benchmarking details and regression equations appear in Supplementary Section S1. PA-equivalent is an interpretive mapping distinct from Relative Change and Coordination Score, which compare observed ALT values to random-policy baselines.
5.4 Random Policy Baselines
For each configuration, we ran 10,000 episodes with uniformly random actions to establish chance-level performance. Coordination is quantified via:
[…]
where $\text{ALT}_{perfect}=1.0$. Negative scores indicate performance worse than random. Random policies achieve deceptively high traditional metrics (Efficiency $\approx 0.82$, Fairness $\approx 0.97$ for 2 agents) but low ALT values (CALT: 0.486 to 0.111 across agent counts).
5.5 Results
Figure 1: Q-Learning vs Random Baseline CALT Performance. Blue bars show Q-learning agents consistently achieving lower CALT values than random baselines (orange bars) across all agent configurations (2, 3, 5, 8, 10 agents). All coordination scores are negative, indicating Q-learning performs worse than chance. Error bars represent standard deviation across Type-A/Type-B and ILF/IQF configurations.
Figure 2: Perfect Alternation Equivalent Performance Degradation. Q-learning agents (blue line; mean across 4 modes) decline from $\sim$56.4% (2 agents) to $\sim$17.9% (10 agents) of perfect coordination, while random baselines (orange line) decline from 69.7% to 33.3%. Shaded regions indicate 95% confidence intervals. Q-learning remains below random across agent counts, indicating persistent coordination deficits.
Figure 3: Traditional Metrics Fail to Detect Coordination Failure. Left panels show traditional outcome metrics achieving moderate-to-high values across configurations, suggesting successful coordination. Right panels show alternation-sensitive metrics exposing poor coordination, with all values below random baselines (dashed lines). This dichotomy demonstrates that traditional outcome-based fairness measures cannot distinguish coordinated alternation from monopolistic or random resource access.
Figure 4: Symmetric Coordination Failure in 3-Agent Systems. Despite nearly identical high traditional metrics (Reward Fairness >> 0.90, Efficiency >> 0.34), both Type-A ILF and Type-B IQF configurations exhibit similarly low ALT performance (CALT $\approx$ 0.14). This symmetric failure across state representations and reward structures demonstrates that coordination deficit is fundamental to independent Q-learning, not an artifact of specific design choices.
Figure 5: CALT Progression During Q-Learning Training. CALT values (solid lines) decrease during learning as epsilon decays (dashed line, right axis), indicating that learned policies achieve worse coordination than early exploration phases. Traditional efficiency (dotted lines) follows a different trend over training and does not track CALT. This finding suggests that convergence to deterministic policies can interfere with alternation while traditional metrics may suggest improvement.
5.5.1 Random Baseline Performance
Table 3 presents comprehensive random baseline results across all agent configurations. The key finding is that random policies achieve deceptively high traditional metrics, creating an illusion of coordination:
Table 3: Random Policy Baseline Performance (10,000 episodes per configuration)
Agents Reward CALT FALT EALT Efficiency Fairness
2 ILF 0.486 0.717 0.651 0.818 0.972
2 IQF 0.486 0.717 0.651 0.818 0.972
3 ILF 0.359 0.630 0.545 0.866 0.967
3 IQF 0.359 0.630 0.545 0.742 0.967
5 ILF 0.243 0.526 0.410 0.727 0.927
5 IQF 0.243 0.526 0.410 0.532 0.927
8 ILF 0.147 0.416 0.251 0.526 0.816
8 IQF 0.147 0.416 0.251 0.305 0.816
10 ILF 0.111 0.363 0.183 0.443 0.846
10 IQF 0.111 0.363 0.183 0.220 0.846
Note: Random policies achieve deceptively high traditional metrics (Efficiency $\geq$ 0.443 for $n\leq 10$, Fairness $\geq$ 0.816), creating an illusion of coordination. In contrast, ALT metrics properly expose lack of coordination, with CALT dropping from 0.486 (2 agents) to 0.111 (10 agents).
For 2-agent systems, random policies achieve Efficiency = 0.818 and Fairness = 0.972—values that would typically be interpreted as indicating successful coordination. Even with 5 agents, random policies maintain Efficiency $\geq 0.532$ and Fairness $\geq 0.927$. These high traditional metric values persist despite the complete absence of any learned coordination strategy.
In stark contrast, ALT metrics properly expose the lack of coordination in random policies. CALT values decrease systematically as agent count increases: from 0.486 (2 agents) to 0.359 (3 agents), 0.243 (5 agents), 0.147 (8 agents), and finally 0.111 (10 agents). This monotonic decrease reflects the increasing difficulty of achieving coordination by chance alone as the number of competing agents grows.
Figure 1 visualizes this critical comparison, showing that Q-learning agents (blue bars) consistently achieve lower CALT values than random baselines (orange bars) across all agent configurations—a finding that challenges assumptions about learned coordination in this class of multi-agent systems. We emphasize that this result does not generalize to all MARL algorithms, but highlights a failure mode of independent tabular learners under temporal coordination demands. We report CALT explicitly as the primary alternation metric; other ALT variants are summarized in Table 4.
5.5.2 Q-Learning Coordination Failure
Our Q-learning experiments reveal that, despite achieving high traditional fairness and efficiency metrics, Q-learning agents demonstrate systematically worse coordination than random policies in this setting. Table 4 reports both Relative Change and Coordination Score for FALT, EALT, CALT, and AALT:
Table 4: Q-Learning vs Random Baseline: Coordination Score Analysis (BASE=1000)
Agents Metric Q-Learning Random Relative Change (%) Coord. Score (%)
2 FALT 0.642 0.717 -10.5% -26.8%
EALT 0.513 0.651 -21.2% -39.6%
CALT 0.322 0.486 -33.7% -31.8%
AALT 0.284 0.435 -34.8% -26.8%
3 FALT 0.480 0.630 -23.8% -40.6%
EALT 0.203 0.545 -62.7% -75.0%
CALT 0.142 0.359 -60.3% -33.7%
AALT 0.140 0.323 -56.5% -26.9%
5 FALT 0.313 0.526 -40.6% -45.0%
EALT 0.099 0.410 -75.9% -52.7%
CALT 0.062 0.243 -74.3% -23.8%
AALT 0.042 0.204 -79.6% -20.4%
8 FALT 0.240 0.416 -42.2% -30.0%
EALT 0.109 0.251 -56.6% -19.0%
CALT 0.047 0.147 -68.2% -11.7%
AALT 0.024 0.103 -76.7% -8.8%
10 FALT 0.232 0.363 -36.2% -20.6%
EALT 0.166 0.183 -9.5% -2.1%
CALT 0.048 0.111 -56.6% -7.1%
AALT 0.022 0.065 -66.6% -4.7%
Key Finding: All coordination scores shown here are negative, indicating Q-learning agents perform up to $\sim$80% worse than random baselines across these ALT metrics. This demonstrates that traditional outcome-based metrics can remain high while temporal coordination is poor.
Relative Change Formula: $\frac{\text{QL}-\text{Random}}{\text{Random}}\times 100$
Coordination Score Formula: $\frac{\text{QL}-\text{Random}}{\text{Perfect}-\text{Random}}\times 100$
Negative values indicate performance below random chance.
Key Finding: For FALT/EALT/CALT/AALT, all Coordination Scores are negative, indicating performance below random baselines after normalization to perfect alternation. Relative Change is negative for all ALT variants except 10-agent qEALT (+5.4%), where more frequent multi-agent ties yield non-zero rewards and behavior drifts toward the random baseline; the quadratic form accentuates small gains and tracks the slightly higher efficiency in that regime (see Supplementary Section S6).
The coordination failure is most severe for specific configurations:
  • 1.
    5 agents, qEALT: -81.2% (worst performance)
  • 2.
    5 agents, AALT: -79.6% (near-worst)
  • 3.
    10 agents, CALT: -56.6% (severe degradation)
Meanwhile, traditional metrics mask this coordination failure completely. For the same experiments, Q-learning agents achieve:
  • 1.
    Reward Fairness: 0.490–0.993 (moderate to near-perfect)
  • 2.
    Efficiency: 0.054–0.677 (low to moderate)
  • 3.
    TT Fairness: 0.699–0.998 (moderate to near-perfect)
These traditional metric values would conventionally be interpreted as successful multi-agent coordination. Our ALT metrics reveal the opposite: Q-learning agents have learned policies that are actively worse than random action selection for achieving turn-taking coordination in this setting.
Figure 3 illustrates this dichotomy, showing high traditional metrics alongside low ALT metrics, demonstrating how conventional measures fail to detect coordination failure in MBoE scenarios.
5.5.3 Perfect Alternation Equivalent Analysis
To provide interpretable quantification of coordination quality, we compute PA-equivalent performance using the AltRatio framework. Table 5 presents this analysis for CALT:
Table 5: Perfect Alternation Equivalent Analysis (CALT - BASE=1000)
Agents Reward CALT ALT Ratio PA Equiv. Performance
Type (QL) Agents (% of Perfect)
2 ILF 0.322 0.568 1.14 56.8%
2 IQF 0.322 0.568 1.14 56.8%
3 ILF 0.142 0.377 1.13 37.7%
3 IQF 0.142 0.377 1.13 37.7%
5 ILF 0.062 0.250 1.25 25.0%
5 IQF 0.062 0.250 1.25 25.0%
8 ILF 0.047 0.216 1.73 21.6%
8 IQF 0.047 0.216 1.73 21.6%
10 ILF 0.048 0.219 2.19 21.9%
10 IQF 0.048 0.219 2.19 21.9%
Interpretation: The ALT Ratio maps observed CALT values to equivalent perfectly alternating agents. For example, 10 Q-learning agents achieve CALT = 0.048, equivalent to only 2.19 perfectly alternating agents (21.9% of perfect coordination).
PA-Equivalent Formula (CALT): $\text{ALT Ratio}=\sqrt{\text{CALT}-1.879\times 10^{-10}}$
$\text{PA Equivalent Agents}=n\times\text{ALT Ratio}$
Key Trend: Performance drops sharply from 2 to 5 agents and remains low for 8–10 agents, demonstrating coordination difficulty scales dramatically with system size.
Table 5 reports reward-type averages, whereas Figure 2 visualizes the mean $\pm$ range across the four modes (Type-A/B $\times$ ILF/IQF).
The PA-equivalent analysis reveals sharp degradation from small to mid agent counts and a low plateau thereafter:
  • 1.
    2 agents: 56.8% of perfect coordination (1.14 equiv. agents)
  • 2.
    3 agents: 37.7% of perfect coordination (1.13 equiv. agents)
  • 3.
    5 agents: 25.0% of perfect coordination (1.25 equiv. agents)
  • 4.
    8 agents: 21.6% of perfect coordination (1.73 equiv. agents)
  • 5.
    10 agents: 21.9% of perfect coordination (2.19 equiv. agents)
Performance drops sharply from 56.8% (2 agents) to 25.0% (5 agents) and remains near $\sim$22% for 8–10 agents, demonstrating that coordination difficulty scales sharply with system size. Notably, 10 Q-learning agents achieve coordination equivalent to only 2.19 perfectly alternating agents—barely one-fifth of the agents in the system exhibiting coordinated behavior.
Figure 2 visualizes this degradation trend, showing how PA-equivalent performance decreases as agent count increases.
5.5.4 Symmetric Coordination Failure
A particularly illuminating case study emerges from comparing 3-agent configurations under different state representations (Figure 4). Despite nearly identical high traditional metrics (Reward Fairness >> 0.90), the systems exhibit similar low ALT performance (CALT $\approx$ 0.14). This symmetric failure across state representations and reward structures suggests that the coordination deficit is fundamental to the Q-learning approach in MBoE, rather than an artifact of specific experimental design choices.
6 Discussion
6.1 Why Does Q-Learning Fail at Coordination?
Our finding—that independent Q-learning agents perform below random baselines—demands explanation. We identify four interrelated factors:
1. The Credit Assignment Problem: Tabular Q-learning cannot recognize that “losing now” (allowing another agent to win) yields “winning later” (reciprocal cooperation). Without function approximation or opponent modeling, agents treat each episode independently, missing the temporal structure of alternation equilibria. The reward for cooperation materializes $n$ episodes in the future, longer than the typical effective temporal credit assignment horizon in tabular Q-learning with $\gamma=0.999$.
2. Non-Stationary Opponent Dynamics: From each agent’s perspective, other agents’ evolving policies violate assumptions underlying standard convergence results for Q-learning. As opponents explore different strategies during early training, the environment appears stochastic and non-stationary. Once all agents simultaneously cease exploration (synchronized epsilon decay at 75% of episodes), they may converge to a suboptimal coordination regime (Nash-like, locally stable) where no single agent can unilaterally improve by alternating—a classic coordination failure Masiliūnas (2019). This synchronized decay may also contribute to mid-range bumps in EALT/qEALT: partial exploration increases exclusive wins before late-stage exploitation drives more simultaneous arrivals and ties.
3. Lack of Explicit Coordination Signals: Unlike humans who achieve PA through implicit signaling (e.g., “I won last time, your turn now”), independent learners have no communication channel. State representations lack sufficient context to infer “whose turn it is,” particularly for $n>3$ agents where alternation cycles exceed the single-episode memory of Type-B states.
4. Tragedy of the Learning Commons: Each agent learns to exploit opportunities for $r_{high}$ greedily, maximizing individual expected return. Without global coordination mechanisms, this individual rationality leads to collective irrationality—agents interfere with each other’s learning trajectories, producing worse outcomes than random policies that at least distribute wins stochastically.
These factors compound as agent count increases, explaining the sharp drop in PA-equivalent performance from 56.8% (2 agents) to 25.0% (5 agents) and the low plateau around $\sim$22% for 8–10 agents. The coordination problem is not merely difficult—it is poorly aligned with the assumptions of independent tabular reinforcement learning.
6.2 Coordination Breakdown as Strategic Regime Shift
The performance drop from 56.8% (2 agents) to $\sim$22% (8–10 agents) reflects a fundamental shift in the equilibrium structure of multi-agent games: as agent count increases, the strategic complexity of achieving coordinated alternation grows sharply, and independent learners fail to discover equilibria supporting fair turn-taking Pikovsky (2009); Maruta and Okada (2012); Masiliūnas (2019); Zhu et al. (2020); Shan et al. (2019); He et al. (2015); Tian et al. (2026). Below a certain agent threshold, residual coordination persists through partial turn-taking patterns; beyond this threshold, individual rationality leads to collective irrationality—agents interfere with each other’s strategies, producing coordination outcomes qualitatively close to random dynamics.
The finding that 10 Q-learning agents coordinate as effectively as only 2.19 perfectly alternating agents constitutes a notable finding, not merely a negative result. It identifies a regime in which coordination breaks down in decentralized multi-agent systems—information that remains entirely hidden when using traditional metrics whose ranges can appear deceptively high (e.g., Reward Fairness up to 0.993 and Efficiency up to 0.677). Such breakdowns, where individual optimization undermines collective welfare, are central concerns in game theory, mechanism design, and the study of social dilemmas.
This perspective reframes our experimental results: rather than demonstrating algorithmic failure, they reveal constraints on emergent coordination in systems where individual optimization conflicts with collective welfare. Performance remains low and close to random between 8 and 10 agents, with only a slight uptick at 10 agents consistent with higher tie frequency and quadratic amplification in qEALT, suggesting a near-random regime that merits finer-grained agent-count sweeps.
6.3 Implications for Random Baseline Methodology
A practical implication of our results is that coordination metrics should be interpreted relative to an explicit null process. Random-policy baselines provide a simple and transparent null hypothesis for “no temporal coordination”: they quantify the level of efficiency, fairness, and alternation that can arise purely from chance under the same environment and episode budget.
In prior work on the Battle of the Exes and related repeated coordination games, random-policy baselines are not consistently reported as a standard reference, and in several cases the presence or absence of such baselines is not discussed explicitly Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023). This can make it difficult to assess whether observed coordination reliably exceeds chance-level behavior, particularly when outcome-based metrics take deceptively high values even under random play.
Additional discussion on interpretability and reconciliation with prior work is provided in Supplementary Section S6.
7 Conclusions
This study introduces Perfect Alternation (PA) as a reference coordination regime and limiting benchmark for temporal turn-taking, alongside six novel ALT metrics and a regression-based benchmarking framework enabling interpretable quantification of coordination quality. Through comprehensive Q-learning experiments with rigorous random policy baselines—a methodological practice not consistently reported in prior BoE-related studies—we reveal fundamental limitations in both traditional evaluation metrics and learned coordination strategies Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023).
7.1 Key Findings
Our experimental results challenge conventional assumptions about multi-agent coordination in three critical ways:
1. Traditional Metrics Severely Fail to Detect Coordination Failure. Across all configurations, Q-learning agents exhibit Reward Fairness in the range 0.49–0.993, Efficiency 0.054–0.677, and TT Fairness 0.699–0.998—values that can appear deceptively high in aggregate. However, ALT metrics reveal the opposite: Relative Change reaches as low as -81% (qEALT), with only one small positive at +5.4% for 10-agent qEALT, demonstrating that learned policies perform worse than random baselines under ALT evaluation. This dichotomy exposes a fundamental inadequacy in outcome-based fairness measures, which cannot distinguish monopolistic, random, or alternating resource access patterns in temporal multi-agent scenarios.
2. Q-Learning Produces Systematic Under-Coordination in Multi-Agent BoE. With only one negligible exception (10-agent qEALT at +5.4%), all coordination scores are negative. This systematic underperformance relative to random baselines (to our knowledge, not previously established in BoE literature) indicates that independent Q-learning agents do not merely fail to achieve coordination—they learn policies that do not promote alternation. The coordination deficit is substantial across sizes (CALT coordination scores: -33.7% for 2 agents, -74.3% for 5 agents, -56.6% for 10 agents), with AALT and qEALT confirming this pattern across the full ALT strictness spectrum, suggesting that tabular Q-learning’s inability to model joint action spaces becomes increasingly problematic as system complexity grows.
3. Coordination Difficulty scales sharply with Agent Count. PA-equivalent analysis shows a steep drop from 56.8% of perfect coordination (2 agents) to 25.0% (5 agents) and a low plateau around $\sim$22% for 8–10 agents. This indicates that coordination challenges compound non-linearly as competitive pressure increases. The finding that 10 agents achieve coordination equivalent to only 2.19 perfectly alternating agents—representing about one-fifth of the system—quantifies the severity of the multi-agent coordination problem in decentralized learning scenarios.
7.2 Theoretical and Methodological Contributions
Perfect Alternation (PA) provides a Pareto-optimal reference point for evaluating turn-taking coordination, analogous to how optimal solutions serve as benchmarks in optimization literature. Unlike irregular or group-based coordination patterns, PA ensures fixed periodicity and temporal fairness regardless of episode count or agent population, making it applicable to indefinite-horizon scenarios common in real-world resource allocation.
ALT Metrics address critical gaps in existing fairness literature by integrating distributional fairness (min/max ratios) with temporal fairness (rotation periodicity). The six metrics offer tunable strictness—from FALT’s lenient fractional evaluation to AALT’s exclusive-win requirement—enabling researchers to select appropriate sensitivity levels for different experimental phases or application domains. Critically, these metrics properly expose coordination failure that traditional measures mask, as demonstrated by our random baseline comparisons.
AltRatio Benchmarking Framework enables interpretable quantification by mapping ALT values to "PA-equivalent agents" through regression equations derived from extreme-case simulations ($n\in[2,40]$). This methodology transforms abstract metric values into actionable statements: "System X coordinates as well as $\sigma$ perfectly alternating agents." The framework generalizes beyond current metrics, providing a foundation for developing measures for hierarchical alternation, weighted distribution, or other coordination regimes.
Random Policy Baseline Methodology provides an explicit statistical null hypothesis for temporal coordination: chance-level behavior under the same environment and evaluation protocol. Because outcome-based metrics can take deceptively high values even under random play, reporting random-policy baselines offers a minimal but crucial reference for interpreting whether coordination genuinely exceeds chance-level performance. In our setting, this reference is decisive: for 10 agents, random policies attain Efficiency in the range 0.22–0.44 and Fairness $\approx 0.846$, while ALT metrics correctly reflect the absence of structured turn-taking.
While random baselines are widely used in other areas of empirical evaluation, they have not been consistently adopted as standard practice in studies of the Battle of the Exes and related coordination games, limiting comparability across works and complicating claims of emergent coordination Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023).
7.3 Implications, Limitations, and Future Directions
Our results highlight constraints of independent tabular learning for temporal coordination and point to promising directions, including richer state representations, explicit communication mechanisms, and centralized training with decentralized execution. Extended implications, limitations, future directions, and application contexts are provided in Supplementary Section S6. We repeated runs with different random seeds and observed qualitatively consistent patterns; for brevity, we report representative results.
By establishing rigorous baselines, developing interpretable benchmarking methodologies, and exposing limitations in conventional evaluation practices, this work advances both theoretical understanding and practical assessment of multi-agent coordination. More broadly, our findings contribute to the social choice and welfare measurement literature by demonstrating that distributional fairness criteria—widely used in fair division and cooperative game theory—can be fundamentally misleading in repeated interaction settings where the temporal sequence of allocations matters as much as their aggregate distribution. The result that Q-learning agents perform below random baselines—despite appearing successful by traditional measures—underscores the necessity of domain-appropriate metrics and the dangers of relying on outcome fairness alone when procedural fairness matters.
Declaration of Generative AI and AI-assisted Technologies in the Writing Process
During preparation of this manuscript, the authors used AI tools (Claude 3.7/4.5; ChatGPT 5.2) for language editing, formatting, and code organization. The authors reviewed and edited all outputs and take full responsibility for the content. The use of these tools did not influence the scientific findings or conclusions.
CRediT authorship contribution statement
Nikolaos Al. Papadopoulos: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data curation, Writing – original draft, Visualization. Konstantinos E. Psannis: Supervision, Writing – review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The code and data supporting the findings of this study are available at https://github.com/dentros/Alternation (GNU GPL). Archived version: https://doi.org/10.5281/zenodo.18528890.
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.